OpenStudy (anonymous):

sqrtx^2-10x+25=25-9x?

OpenStudy (anonymous):

\[\sqrt{x ^{2}-10x+25}\]=25-9x?

OpenStudy (anonymous):

\[x^2-10x+25=(x-5)^2\] is as start

OpenStudy (anonymous):

(x-5) is that part of the answer